OpenStudy (anonymous):

Factor the polynomial completely. 6x^2 - 19x + 13

OpenStudy (anonymous):

i hate factoring

OpenStudy (anonymous):

try \((x-1)(6x-13)\)

OpenStudy (campbell_st):

well start by multiplying 6 by 13 6 * 13 = 78 find the factors of 78 that add to - 19 they are both negative... one of then is -6, whats the other...?

OpenStudy (dumbcow):

multiply outside coefficients ..... 6*13 = 78 now try to find 2 factors of 78 that add up to -19 -6 + -13 = -19 -6 * -19 = 78 now once you have found 2 factors that work, you can use factor by grouping or there is a shortcut method bring down the first coefficient, then reduce any common factors \[\rightarrow (6x-6)(6x-13)\] \[\rightarrow (x-1)(6x-13)\]

OpenStudy (dumbcow):

sorry @campbell_st , didn't see your post

OpenStudy (campbell_st):

thats fine...@dumbcow just so people can see what you cancel I always include the steps (6x - 6)(6x - 13) 6(x -1)(6x -13) -------------- = ---------------- 6 6 cancel the common factor.